OpenStudy (anonymous):

Two numbers have a sum of 18. If the small number doubled, and then subtracted from the larger number, the difference is 9. X represents the smaller number, Y represents the biggest number. Write an equation to model the difference of the numbers.

OpenStudy (anonymous):

let write the system of equatios let L be the larger one and s for smaller one their sum is 18 s+L=18 and the second statement says: L-2*s=9 so s+L=18 -2s+L=9 Ca you make it to end now?